OpenStudy (anonymous):

ABCD is a rectangle A is the point= (0,1) C is the point= (0,5) the equation of the straight line through A and B is y= 2x+1 (a) find the equation of the straight line through d and C @Godlovesme

OpenStudy (anonymous):

so, how do i start this

OpenStudy (godlovesme):

this is confusing lol i'm trying to figure it out :/

OpenStudy (anonymous):

ok, i agree

OpenStudy (jdoe0001):

is there a pcture?

OpenStudy (anonymous):

np

OpenStudy (anonymous):

so, what do you think

OpenStudy (anonymous):



OpenStudy (anonymous):

this is the diagram. by the way there is a C on the right side of 6

OpenStudy (dtan5457):

Ok. I got it

OpenStudy (anonymous):

yes, pls share

OpenStudy (dtan5457):

When you plot these points, and the line. You are given the information that a=0,1 c=0,5 b=2,5 d=2,1 to make an equation that passes through both 0,5 and 2,1 (c and d) get the slope, which is -2x point slope= y-1=-2(x-2) y=-2x+5

OpenStudy (dtan5457):

And when you go ahead and graph that line, it goes straight through c and D

OpenStudy (anonymous):

thank you soo much. now for part 2 what is the equation of the straight line through B and C

OpenStudy (dtan5457):

(2,5)(2,1)<<<b and C a vertical line would go through that x=?

OpenStudy (jdoe0001):

notice your graph bear in mind that a rectangle has all right-angles, so all 4 angles are 90 degrees each  so